OpenStudy (anonymous):

Help please! I forgot how to do this. If f(x) = 0.8(2 – x), what is the value of f(–2)? A. 0 B. 1.6 C. 3.2 D. 3.6

OpenStudy (anonymous):

@ranga

OpenStudy (ranga):

Substitute x = -2 and compute the right hand side. f(x) = 0.8(2 – x) f(-2) = 0.8(2 - (-2)) = ?

OpenStudy (anonymous):

3.2?

OpenStudy (ranga):

Yes.

OpenStudy (anonymous):

that's it?

geerky42 (geerky42):

Yes.

OpenStudy (anonymous):

thanks! both of u!